OpenStudy (anonymous):

You’ve been working the box office at a movie theater for a year now and are anxious for a promotion. Your supervisor asks you to find out how many adults and how many children attended the recent screening of “Amazing Algebra 2.” The total number of adult and child tickets sold was 150. Each adult ticket costs $9 and one child ticket costs $6. The movie theater made $1,155. Looks like you’ll need some Algebra to solve this one! Start by identifying the variables and writing the first equation. You’re counting the number of adult (a) tickets and the number of children (c) tickets....

OpenStudy (anonymous):

The total number sold is 150 tickets. So, what do you think the first equation will look like?

OpenStudy (anonymous):

\[a+c=150\]

OpenStudy (anonymous):

if \(a\) is the number of adult tickets solve and \(c\) is the number of children's tickets, then to total is \(a+b\) and you know the total is 150 so one equation is \(a+b=150\)

OpenStudy (anonymous):

and to anticipate the next question, \(a\) adult tickets bring a total of \(9a\) dollars and \(c\) child's tickets earn \(6c\) dollars so \[9a+6c=1155\]

OpenStudy (ganpat):

1st eqn would be, total no of tickets sold.. x +y y = 150; 2nd eqnt will be, no of adults and no of kids.. 9x + 6y = 1155.. so by solving them simultaneously, we get.. x = 85 and y = 65...

OpenStudy (anonymous):

Now, the second equation needs to be written. Adult tickets cost $9, so the total earned from the number of adult tickets sold (a) is represented by the term 9a. Child tickets cost $6, so the total earned from child tickets (c) is represented by the term 6c. You know the movie earned $1,155. Can you figure out the second equation? Type it in the space provided. (No dollar signs needed.)

OpenStudy (ganpat):

9a + 6c = 1155.. is the 2nd equation...